OpenStudy (anonymous):

explain how is (2, -4) an ordered pair for this equation 2x + 3y = -8

OpenStudy (anonymous):

An ordered pair is of the form (x,y), so let's plus 2 in for x and -4 in for y and see if it equals -8. 2*2 + 3*-8 = 4-12 = -8.

OpenStudy (anonymous):

since there is already a given ordered pair, all you have to do is to check whether the given ordered pair satisfies the equation. step: substitute the value of x and y, given the ordered pairs and solve if it's equal to -8. it it's equal then it's an ordered pair for the equation. ..